OpenStudy (anonymous):

The following are the slopes of the lines repre- senting daily revenues y in terms of time x in days. Use the slopes to interpret any change in daily revenues for a 1-day increase in time. a) m = 400 1. The revenues are decreasing $400 per day 2. Unable to determine 3. The revenues are increasing $400 per day

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Why not explore it anc create your own experiment? y = 400x + b Try something. Maybe x = 10. Okay, now try x = 11 - a one day increase. What happened?

OpenStudy (anonymous):

the revenues are decreasing 400 per day?

OpenStudy (tkhunny):

Notice how one need not find the intercept at all!

OpenStudy (anonymous):

@lasttccasey can you help me

OpenStudy (lasttccasey):

Well a positive slope generally means increase over the x interval which in this case would be time. So I would assume the answer to be: 3. The revenues are increasing $400 per day

OpenStudy (anonymous):

@raffle_snaffle

OpenStudy (tkhunny):

Did you compare x = 10 and x = 11? It will tell you what happens. y = 400x + b

OpenStudy (anonymous):

let me do it hang on

OpenStudy (anonymous):

increase?

OpenStudy (tkhunny):

(400(11) + b) - (400(10) + b) = 400(11-10) = 400 Oh, so the increase of one day creates an increase of 400.

OpenStudy (anonymous):

@tkhunny how about m=100?

OpenStudy (tkhunny):

Explore it!! Do what I just did and see what happens.

OpenStudy (anonymous):

@tkhunny it would be the same thing, an increase right

OpenStudy (tkhunny):

Slope is 100, increase is 100. Slope is 400, increase is 400. Slope is -40, decrease is 40. It's the nature of a linear function. It all happens together and at the same rate and in the same direction.

OpenStudy (anonymous):

how about if it was 0?

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Why would it be any different?